OpenStudy (anonymous):

Critical Points? \[(2x-5^{3})(x+4^{2})\]

OpenStudy (anonymous):

Did you take the derivative?

OpenStudy (anonymous):

\[f \prime (x)= 6(2x-5)^{2} + (x+4) ^{2} + 2 (2x-5)^{3} (x+4)\]

OpenStudy (amistre64):

5^3 and 4^2 are constants did you account for that?

OpenStudy (amistre64):

f'(x) = 2(x+4^2) + (2x-5^3)(1)

OpenStudy (anonymous):

\[(2x-5^3)(x+4^2)\] \[=2x^2 + 2(4^2)x - 5^3x -20\] \[\frac{d}{dx}[2x^2 + 8x - 5^3x -20]\] \[= 4x + 2(4^2) - 5^3 - 0\] \[=4x + 32 - 125\] \[= 4x - 93\]